OpenStudy (anonymous):

7.8y+9=6.8y-2 how do i solve this?

OpenStudy (kainui):

Since it has an equals sign, it means that the left is equal to the right. So what you can do is add or subtract something to both sides, and it stays the same, right? So in this case we can subtract 6.8y from both sides and get y+9=-2 Now do it one more time to get y by itself.

OpenStudy (anonymous):

7.8y+9=6.8y-2 56y+9=48y-2 56y-48y=-2-9 8y=-11 y= -11/8

OpenStudy (kainui):

@shubham.bagrecha That's just wrong.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@mf0000 7.8 means 7*8 or its decimal point?

OpenStudy (anonymous):

*

OpenStudy (anonymous):

@shubham.bagrecha Are you @mf0000 ?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

No its a decimal point.

OpenStudy (anonymous):

@Kainui check yours. it will be 8y= -11.

OpenStudy (anonymous):

@curiousshubham just use common sense.

OpenStudy (anonymous):

7.8y + 9 = 6.8y - 2 y = -2 - 9 = -11

OpenStudy (kainui):

Common sense tells me that it's a decimal point. Decimal points exist in mathematics lol.

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

decimal...

OpenStudy (kainui):

mf, just listen to what my first response was. pratu has the right answer, he finished the problem for you.

OpenStudy (anonymous):

@shubham.bagrecha don't use your common sense too much....

OpenStudy (anonymous):

yeah @pratu043 its -11.

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

welcome and welcome to OS

OpenStudy (anonymous):

7x+9=6x-2 , x= 8y 7x-6x = -2-9 x= -11 8y=-11